# Brain Organoid Computing – an Overview

### v1

@@ V Discussion
Human brain organoids have been shown to be capable of highly efficient, parallelized computation, with extrapolation capacities that ANNs lack, all while current experiments are rather low-cost. Nevertheless, these capabilities are hindered by above mentioned challenges. The development and application of human brain organoids is still in its formative stages, with current in vitro cultures still being rather limited. These issues are compounded by challenges of reliably reproducing results, due to the heterogeneous nature of brain organoids. Furthermore, ethical and legal implications of experimenting on and using human brain cells as a tool for computations necessitate close monitoring and work in accordance with current regulations for further research to be possible.
@@
[…]
[Uncaptioned image] Yannic Talavera was born in Worms, Germany in 2002. He completed his bachelor’s degree in Business Informatics at the FOM University of Applied Sciences in 2024. His academic interests include unconventional computing approaches, cognitive science, and their interdisciplinary applications.
[Uncaptioned image] Bernd Ulmann was born in Neu-Ulm, Germany in 1970, and received his diploma in mathematics from the Johannes Gutenberg-Universität Mainz, Germany, in 1996. He received his Ph.D. from the Universität Hamburg, Germany, in 2009. Since 2010 he is professor for business informatics at the FOM University of Applied Sciences, Frankfurt/Main, Germany. His main interests are analog and hybrid computing, the simulation of dynamic systems and operator methods. He is author of several books on analog and hybrid computing.

### v2

@@ V Discussion
Human brain organoids have been shown to be capable of highly efficient, parallelized computation, with extrapolation capacities that ANNs lack, all while current experiments are rather low-cost. Nevertheless, these capabilities are hindered by above-mentioned challenges. The development and application of human brain organoids is still in its formative stages, with current in vitro cultures still being rather limited. These issues are compounded by challenges of reliably reproducing results, due to the heterogeneous nature of brain organoids. Furthermore, ethical and legal implications of experimenting on and using human brain cells as a tool for computations necessitate close monitoring and work in accordance with current regulations for further research to be possible.
@@
In the end, if the biological complexities of these systems are solved, human brain organoid computing could become a new frontier for multidisciplinary computing technologies. For the time being, however, researchers must focus on fundamental research to build on existing findings and overcome existing challenges to move forward to real-world applications. The full potential of brain organoid technology in computing is yet to be realized, offering a range of possibilities to be explored in future research.